OpenStudy (anonymous):

Find the least common multiple of each pair of polynomials. 3x+5 and 9x^2-25

OpenStudy (helder_edwin):

did u factor the quadratic polynomial. do it. it will give u a clue

OpenStudy (helder_edwin):

u can factor \[\large 9x^2-25=(3x)^2-5^2=(3x+5)(3x-5) \]

OpenStudy (helder_edwin):

so \[\large \text{l.c.m.}(3x+5;9x^2-25)=(3x+5)(3x-5)=9x^2-25\]

OpenStudy (helder_edwin):

got it?

OpenStudy (anonymous):

yes, thank you!!

OpenStudy (helder_edwin):

u r welcome